# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 5 WEF Solutions for binary additive Valuations
Assume $\ell_{i}(A^{t})=(i=i_{1},...,i_{k})$ is the path with the highest total cost starting at $i$ in the weighted envy graph corresponding to the allocation at the end of the iteration, $G_{A^{t},w}$.
Observation 5.6 implies $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{|A_{i_{k}}^{t}|}{w_{i_{k}}}-\frac{|A_{i}^{t}|}{w_{i}}$.
As $i\in U^{t}$, Lemma 5.9 implies $\frac{|A_{i_{k}}^{t}|}{w_{i_{k}}}-\frac{|A_{i}^{t}|}{w_{i}}\leq\frac{1}{w_{i}}$. ∎
Proposition 5.10 ensures an upper bound on the subsidy required for agents remaining in the game: for each agent $i\in U^{t}$ and iteration $t\in T$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{i}}\leq\frac{1}{w_{1}}$.
Note that, in the special case of identical binary valuations, all agents remain in the game simultaneously. Therefore, 5.10 ensures an upper bound of $w_{i}\cdot\frac{1}{w_{i}}=1$ on the subsidy of each agent, which implies a total subsidy of at most $n-1$. However, this subsidy bound is not applicable for agents who were removed from the game. For an agent $i$ removed from the game, we can prove a subsidy bound of $w_{i}\cdot\frac{1}{w_{j}}$, where $j$ is some other agent that remains in the game.
To prove this upper bound, we need to establish several claims about agents who are removed from the game. First, we show that an agent removed from the game does not desire any item held by an agent who remains in the game (5.11). As a result, these removed agents will not be included in any transfer path (5.12).
Next, we demonstrate that if the cost of a path originating from these removed agents at some iteration $t$ exceeds the cost at $t^{\prime}\leq t$ — the iteration when the agent was removed — then there exists an arc in this path, $(i_{j},i_{j+1})$, such that $v_{i}(A_{j}^{t})=0$ (5.13). Based on these claims, we prove that if at iteration $t$, the cost of the maximum-cost path starting from agent removed from the game at $t^{\prime}<t$ exceeds its cost at $t^{\prime}$, we can upper-bound it by $\frac{1}{w_{1}}$.
Proposition 5.11.
Let $i$ be an agent removed from the game in Phase 1 at iteration $t^{\prime}$. Then for all $j\in U^{t^{\prime}}$ who were not removed in Phase 1 at iteration $t^{\prime}$, $v_{i}(A^{t^{\prime}}_{j})=0$.
Moreover, for all $t>t^{\prime}$ and all $j\in U^{t}$, $v_{i}(A^{t}_{j})=0$.
@@
Suppose that, at iteration $t^{\prime}$, after Stage 2, $v_{i}(A_{j}^{t^{\prime}})\neq 0$. This implies there exists some item $o\in A_{j}^{t^{\prime}}$ such that $v_{i}(o)=1$. We consider two cases:
@@
    $o\in A_{j}^{t^{\prime}-1}$. In this case, at the beginning of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j$. Moreover, there is a transfer path from $j$ to $i_{0}$ at iteration $t^{\prime}$ (otherwise, $j$ would have been removed from the game at $t^{\prime}$ as well). Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
@@
    $o\not\in A_{j}^{t^{\prime}-1}$, that is, $j$ received item $o$ during iteration $t^{\prime}$, from some other agent $j^{\prime}$ (where $j^{\prime}=i_{0}$ is possible). At the beginning of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j^{\prime}$. Moreover, there is a transfer path from $j^{\prime}$ to $i_{0}$, which is used to transfer the newly allocated item. Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
Both cases contradict the assumption that $i$ removed at $t^{\prime}$.
@@
We assume the claim holds for iteration $t-1>t^{\prime}$ and prove it for iteration $t$. Assume, contrary to the claim, that there exists an agent $i$ who was removed at iteration $t^{\prime}$, and an agent $j\in U^{t}$ such that $v_{i}(A_{j}^{t})\neq 0$. By the induction hypothesis and the fact that $U^{t-1}\supseteq U^{t}$, we have $v_{i}(A_{j}^{t-1})=0$. Therefore, during iteration $t$, $j$ must have received a new item $o_{j}$ that $i$ values at $1$.
If $o_{j}$ was part of $i_{0}$’s bundle at $t^{\prime}$, then the transfer path starting at $i$, $(i,i_{0})$, must have already existed at $t^{\prime}$.
Alternatively, if $o_{j}$ was originally in another agent’s bundle at $t^{\prime}$, say agent $k\in N$, then there must have been an iteration between $t^{\prime}$ and $t$ in which $o_{j}$ has been transferred from $k$ to another agent, while agent $k$ is compensated by some other item $o_{k}$, that agent $k$ wants.
If $o_{k}$ was part of $i_{0}$’s bundle at $t^{\prime}$, then the transfer path starting at $i$, $(i,k,i_{0})$, must have existed at $t^{\prime}$. Otherwise, $o_{k}$ was in another agent’s bundle at $t^{\prime}$, and it, too, would be transferred to a different bundle in a later iteration.
Since the number of items is finite, this process must eventually lead to an item that was in $A_{i_{0}}^{t^{\prime}}$, forming a transfer path starting at $i$ at iteration $t^{\prime}$ — a contradiction.∎
Proposition 5.12.
Let $i$ be an agent removed from the game in Stage 1 at iteration $t^{\prime}$. Then, for all $t\geq t^{\prime}$, $i$ will not be included in any transfer path.
@@
First, note that $v_{i}(A_{i_{0}}^{t})=0$; otherwise, $i$ would not have been removed at $t^{\prime}$. Next, any agent $j$, who receives an item from $i_{0}$’s bundle in iteration $t$, must be in $U^{t}$. By Proposition 5.11, $v_{i}(A_{j}^{t})=0$, which means that $i$ can not receive any item from $j\in U^{t}$ in compensation for another item from their own bundle. Thus, any transfer path includes only agents in $U^{t}$. ∎
Proposition 5.13.
Consider an iteration $t$ and an agent $i\notin U^{t}$, who was removed from the game in Phase 1 at iteration $t^{\prime}<t$. Let $L=(i=i_{1},\ldots,i_{k})$ be a path in the weighted envy graph starting at $i$. If $cost_{A^{t}}(L)>cost_{A^{t^{\prime}-1}}(L)$, then there must exist $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin U^{t}$, $i_{j+1}\in U^{t}$ and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$.
@@
Let $t>t^{\prime}$ be the earliest iteration in which $cost_{A^{t}}(L)>cost_{A^{t^{\prime}-1}}(L)$. This implies that there is at least one agent, say $11\leq j^{\prime}\leq k-$ 1, such that agent $i_{j^{\prime}+1}$ has received a new item that $i_{j^{\prime}}$ desires. In other words, $i_{j^{\prime}+1}$ was part of a transfer path at iteration $t$, and by Proposition 5.12, $i_{j^{\prime}+1}\in U^{t}$ (in particular, $j^{\prime}\geq 2$).
Since $i_{1}\notin U^{t}$ and $i_{j^{\prime}+1}\in U^{t}$, there must exist some $11\leq p<j^{\prime}+$ 1 such that $i_{p}\notin U^{t}$ and $i_{p+1}\in U^{t}$. By Proposition 5.11, $v_{i_{p}}(A^{t}_{i_{p+1}})=0$. ∎
Proposition 5.14.
Let $i\notin U^{t}$, be an agent who was removed from the game at iteration $t^{\prime}<t$. Then, at the end of iteration $t$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{1}}$.
@@
Recall that $\ell_{i}(A^{t^{\prime}-1})$ is the highest-cost path starting from $i$ at iteration $t^{\prime}-1$ (before agent $i$ removed). In particular, $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))\leq cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))$. Moreover, by Observation 5.6 we have $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))\leq\frac{|A_{i_{k}}^{t^{\prime}-1}|}{w_{i_{k}}}-\frac{|A_{i}^{t^{\prime}-1}|}{w_{i}}$ when $i_{k}$ is the last agent in $\ell_{i}(A^{t^{\prime}-1})$. Combined with Lemma 5.9, this gives $cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t^{\prime}-1}))\leq\frac{1}{w_{i}}$. Therefore, if $cost_{A^{t}}(\ell_{i}(A^{t}))\leq cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))$, then we are done.
Assume now that $cost_{A^{t}}(\ell_{i}(A^{t}))>cost_{A^{t^{\prime}-1}}(\ell_{i}(A^{t}))$. Denote the path $\ell_{i}(A^{t})$ by $(i=i_{1},\ldots,i_{k})$.
From Proposition 5.13, there exists $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin U^{t}$, $i_{j+1}\in U^{t}$ and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$. Then, by Observation 5.6:
$cost_{A^{t}}(\ell_{i}(A^{t}))=$ (4)
$cost_{A^{t}}(i,...,i_{j})+cost_{A^{t}}(i_{j},i_{j+1})+cost_{A^{t}}(i_{j+1},...,i_{k})\leq$
$\leq\left(\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}-\frac{|A_{i}^{t}|}{w_{i}}\right)+\left(0-\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}\right)+cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))\leq$
$cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))$ (5)
Since $i_{j+1}\in U^{t}$, it follows from (5) and Proposition 5.10 that
$cost_{A^{t}}(\ell_{i}(A^{t}))\leq cost_{A^{t}}(\ell_{i_{j+1}}(A^{t}))\leq\frac{1}{w_{i_{j+1}}}\leq\frac{1}{w_{1}}.$
We are now prepared to establish the proof of Theorem 5.8.
@@
Together Proposition 5.10 and Proposition 5.14 establish that for every $i\in N$ and $t\in[T]$, $cost_{A^{t}}(\ell_{i}(A^{t}))\leq\frac{1}{w_{1}}$. Along with Lemma 5.7 and Lemma 5.5, Algorithm 3 computes a WEF-able allocation $A^{T}$ where the required subsidy per agent $i\in N$ is at most $w_{i}\cdot\frac{1}{w_{1}}=\frac{w_{i}}{w_{1}}$.
As there is at least one agent who requires no subsidy (see Theorem 3.4), the total required subsidy is at most $\frac{W-w_{1}}{w_{1}}=\frac{W}{w_{1}}-1$.
We complete the proof of Theorem 5.8 by demonstrating that Algorithm 3 runs in polynomial-time. We represent the valuations using a binary matrix $X$ where $v_{i}(o_{j})=1\Longleftrightarrow X(i,j)=1$. Hence, the allocation of items to the bundle of $i_{0}$ at line 1 can be accomplished in $O(mn)$ time.
At each iteration $t$ of the while loop, either $A_{i_{0}}$ or $U^{t}$ reduced by 1, ensuring that the loop runs at most $m+n$ times.
@@
According to Viswanathan et al. [26], finding a transfer path starting from agent $i\in N$ (or determining that no such path exists) takes $O(T_{v}\log m)$. Removing agents at the start of each iteration incurs a complexity of $O(nT_{v}\log m)$. Furthermore, as stated in Viswanathan et al. [26], identifying $i^{t}$ requires $O(nT_{v})$. Updating the allocation based on the transfer path, according to the same source, takes $O(m)$.
Thus, each iteration has a total complexity of $O(nT_{v}\log m+nT_{v}+T_{v}\log m+m)=O\left(nT_{v}\log m+m\right)$.
@@
A discussion about the tightness of the bound can be found in appendix A.2
Notice that since the output allocation from Algorithm 3 is non-redundant, $A$ maximizes the USW . Moreover, Example 5.3, is not WEF(1,0) (No matter which item is removed from $i_{2}$’s bundle, $i_{1}$ will still feel envious). Hence we can conclude that a $MUSW^{M,N,v}$ allocation is not necessarily WEF(1,0) either. However, it is WEF(0,1).
Proposition 5.15.
For additive binary valuations, Algorithm 3 computes a $WEF(0,1)$ allocation.
@@
We prove by induction that at each iteration $t\in[T]$ $A^{t}$ satisfies $WEF(0,1)$. This means that for every $i,j\in N$, there exists a set of items $B\subseteq A_{j}^{t}$ of size at most $1$ such that $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}$.
@@
    $v_{i}(A_{j}^{t})=v_{i}(A_{j}^{t-1})$: This is the case where $j$ was not included in a transfer path, or was included but was not the first agent in the path, and exchanged an item for a new one, both having the same value for $i$. By the induction assumption, there exists some singleton $B^{t-1}\subseteq A_{j}^{t-1}$ such that $\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$. There exists some singleton $B^{t}\subseteq A_{j}^{t}$, with $v_{i}(B^{t})=v_{i}(B^{t-1})$. Hence, $\frac{v_{i}(A_{i}^{t})+v_{i}(B^{t})}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$.
@@
    $v_{i}(A_{j}^{t})<v_{i}(A_{j}^{t-1})$: This is the case where $j$ was included in a transfer path but exchanged an item $i$ values for an item $i$ does not value. Then $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}>\frac{v_{i}(A_{j}^{t})}{w_{j}}$ for a set $B\subseteq A_{j}^{t}$ of size at most $1$.
@@
      We can conclude that $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+1}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})+1}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$. The claim holds for $B=\{o\}\subseteq A_{j}^{t}$.
@@
      $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}<\frac{v_{i}(A_{j}^{t})}{w_{j}}\leq\frac{v_{j}(A_{j}^{t})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq$
      $\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}\leq\frac{v_{i}(A_{i}^{t})+1}{w_{i}},$

### v4

@@ 5 WEF Solutions for binary additive Valuations
Assume $P_{i}=(i,\ldots,j)$ is the path with the highest total cost starting at $i$ in the $G_{A^{t},w}$, i.e., $cost_{A^{t}}(P_{i})=\ell_{i}(A^{t})$. Section 5 implies $\ell_{i}(A^{t})\leq\frac{|A_{j}^{t}|}{w_{j}}-\frac{|A_{i}^{t}|}{w_{i}}$. As $i\in R(t)$, Section 5 implies $\frac{|A_{j}^{t}|}{w_{j}}-\frac{|A_{i}^{t}|}{w_{i}}\leq\frac{1}{w_{i}}$. ∎
To prove this upper bound, we need to establish several claims about agents removed from the game. First, we show that an agent removed from the game does not desire any item held by an agent who remains in the game (Section 5). As a result, these removed agents will not be included in any transfer path (Section 5).
Next, we demonstrate that if the cost of a path originating from one of these removed agents at the end of iteration $t$ exceeds the cost at the end of iteration $t^{\prime}\leq t$ — the iteration when the agent was removed — then there exists an edge in this path, $(i_{j},i_{j+1})$, such that $v_{i}(A_{j}^{t})=0$ (Section 5). Based on these claims, we prove that if at the end of iteration $t$, the cost of the maximum-cost path starting from agent removed from the game at $t^{\prime}<t$ exceeds its cost at $t^{\prime}$, we can upper-bound it by $\frac{1}{w_{1}}$.
{proposition}
Let $i$ be an agent removed from the game at the start of iteration $t^{\prime}$. Then for all $j\in R(t^{\prime})$, $v_{i}(A^{t^{\prime}}_{j})=0$.
Moreover, for all $t>t^{\prime}$ and all $j\in R(t)$, $v_{i}(A^{t}_{j})=0$.
@@
Suppose that $v_{i}(A_{j}^{t^{\prime}})\neq 0$. This implies there exists some item $o\in A_{j}^{t^{\prime}}$ such that $v_{i}(o)=1$. We consider two cases:
@@
    $o\in A_{j}^{t^{\prime}-1}$. In this case, at the start of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j$. Moreover, there is a transfer path from $j$ to $i_{0}$ at the start of iteration $t^{\prime}$ (otherwise, $j$ would have been removed from the game at $t^{\prime}$ as well). Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
@@
    $o\not\in A_{j}^{t^{\prime}-1}$, that is, $j$ received item $o$ during iteration $t^{\prime}$, from some other agent $j^{\prime}$ (where $j^{\prime}=i_{0}$ is possible). At the start of iteration $t^{\prime}$, there exists a transfer path from $i$ to $j^{\prime}$. Moreover, there is a transfer path from $j^{\prime}$ to $i_{0}$, which is used to transfer the newly allocated item. Concatenating these paths gives a transfer path from $i$ to $i_{0}$.
Both cases contradict the assumption that $i$ was removed at $t^{\prime}$.
@@
We assume the claim holds for iteration $t-1>t^{\prime}$ and prove it for iteration $t$. Assume, contrary to the claim, that there exists an agent $i$ who was removed at the start of iteration $t^{\prime}$, and an agent $j\in R(t)$, such that $v_{i}(A_{j}^{t})\neq 0$. By the induction hypothesis and the fact that an agent can not be added to $R$, we have $v_{i}(A_{j}^{t-1})=0$. Therefore, during iteration $t$, $j$ must have received a new item $o_{j}$ that $i$ values at $1$.
If $o_{j}$ was part of $i_{0}$’s bundle at the start of iteration $t^{\prime}$, then the transfer path starting at $i$, $(i,i_{0})$, must have already existed at the start of iteration $t^{\prime}$.
Alternatively, if $o_{j}$ was originally in another agent’s bundle at the start of iteration $t^{\prime}$, say agent $k\in N$, then there must have been an iteration between $t^{\prime}$ and $t$ in which $o_{j}$ has been transferred from $k$ to another agent, while agent $k$ is compensated by some other item $o_{k}$, that agent $k$ wants.
If $o_{k}$ was part of $i_{0}$’s bundle at the start of iteration $t^{\prime}$, then the transfer path starting at $i$, $(i,k,i_{0})$, must have existed at the start of iteration $t^{\prime}$. Otherwise, $o_{k}$ was in another agent’s bundle at $t^{\prime}$, and it, too, would be transferred to a different bundle in a later iteration.
Since the number of items is finite, this process must eventually lead to an item that was in $A_{i_{0}}^{t^{\prime}}$, forming a transfer path starting at $i$ at the start of iteration $t^{\prime}$ — a contradiction. ∎
{proposition}
Let $i$ be an agent removed from the game at the start of iteration $t^{\prime}$. Then, for all $t\geq t^{\prime}$, $i$ will not be included in any transfer path.
@@
First, note that $v_{i}(A_{i_{0}}^{t})=0$; otherwise, $i$ would not have been removed at the start of iteration $t^{\prime}$. Next, any agent $j$, who receives an item from $i_{0}$’s bundle at iteration $t$, must be in $R(t)$. By Section 5, $v_{i}(A_{j}^{t})=0$, which means that $i$ can not receive any item from $j$, who in $R(t)$, as compensation for another item from their own bundle. Thus, any transfer path in iteration $t$ includes only agents in $R(t)$. ∎
{proposition}
Consider an iteration $t$ and an agent $i\notin R(t)$, who was removed from the game at the start of iteration $t^{\prime}<t$. Let $P=(i=i_{1},\ldots,i_{k})$ be a path in the weighted envy graph starting at $i$. If $cost_{A^{t}}(P)>cost_{A^{t^{\prime}-1}}(P)$, then there must exist $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin R$ at $t$, $i_{j+1}\in R(t)$ at $t$, and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$.
@@
Let $t>t^{\prime}$ be the earliest iteration in which $cost_{A^{t}}(P)>cost_{A^{t^{\prime}-1}}(P)$. This implies that there is at least one agent, say $11\leq j^{\prime}\leq k-$ 1, such that agent $i_{j^{\prime}+1}$ has received a new item that $i_{j^{\prime}}$ desires. In other words, $i_{j^{\prime}+1}$ was part of a transfer path at the start of iteration $t$, and by Section 5, $i_{j^{\prime}+1}\in R(t)$ (in particular, $j^{\prime}\geq 2$).
Since $i_{1}\notin R(t)$ and $i_{j^{\prime}+1}\in R(t)$, there must exist some $11\leq p<j^{\prime}+$ 1 such that $i_{p}\notin R(t)$ at $t$ and $i_{p+1}\in R(t)$. By Section 5, $v_{i_{p}}(A^{t}_{i_{p+1}})=0$. ∎
{proposition}
Let $i\notin R(t)$, be an agent who was removed from the game at the start of iteration $t^{\prime}<t$. Then, for $A^{t}$ the resulting allocation from iteration $t$, $\ell_{i}(A^{t})\leq\frac{1}{w_{1}}$.
@@
Denote by $P_{i}^{t^{\prime}-1},P_{i}^{t}$ the highest-cost paths starting from $i$ at iterations $t^{\prime}-1$ (before agent $i$ removed) and $t$, correspondingly. In particular,
$cost_{A^{t^{\prime}-1}}(P_{i}^{t})\leq cost_{A^{t^{\prime}-1}}(P_{i}^{t^{\prime}-1})=\ell_{i}(A^{t^{\prime}-1}).$
Moreover, by Section 5 we have
$cost_{A^{t^{\prime}-1}}(P_{i}^{t^{\prime}-1})\leq\frac{|A_{i_{k}}^{t^{\prime}-1}|}{w_{i_{k}}}-\frac{|A_{i}^{t^{\prime}-1}|}{w_{i}}$
when $i_{k}$ is the last agent in $P_{i}^{t^{\prime}-1}$. Combined with Section 5, this gives $\ell_{i}(A^{t^{\prime}-1})\leq\frac{1}{w_{i}}$. Therefore, if $cost_{A^{t}}(P_{i}^{t})=\ell_{i}(A^{t})\leq\ell_{i}(A^{t^{\prime}-1})$, then we are done.
Assume now that $\ell_{i}(A^{t})>\ell_{i}(A^{t^{\prime}-1})$. Denote the path $P_{i}^{t}$ by $(i=i_{1},\ldots,i_{k})$.
From Section 5, there exists $j\in\{1,\ldots,k-1\}$ such that $i_{j}\notin R(t)$, $i_{j+1}\in R(t)$ and $v_{i_{j}}(A_{i_{j+1}}^{t})=0$. Then, by Section 5:
$\ell_{i}(A^{t})=cost_{A^{t}}(i,...,i_{j})+cost_{A^{t}}(i_{j},i_{j+1})+cost_{A^{t}}(i_{j+1},...,i_{k})\leq$
$\leq\left(\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}-\frac{|A_{i}^{t}|}{w_{i}}\right)+\left(0-\frac{|A_{i_{j}}^{t}|}{w_{i_{j}}}\right)+\ell_{i_{j+1}}(A^{t})\leq$
$\ell_{i_{j+1}}(A^{t}).$ (4)
Since $i_{j+1}\in R(t)$, it follows from (4) and Section 5 that
$\ell_{i}(A^{t})\leq\ell_{i_{j+1}}(A^{t})\leq\frac{1}{w_{i_{j+1}}}\leq\frac{1}{w_{1}}.$
{theorem}
@@
Together Section 5 and Section 5 establish that for every $i\in N$ and $t\in[T]$, $\ell_{i}(A^{t})\leq\frac{1}{w_{1}}$. Along with Algorithm 3, Algorithm 3 computes a WEF-able allocation $A^{T}$ where the required subsidy per agent $i\in N$ is at most $\frac{w_{i}}{w_{1}}$.
As there is at least one agent who requires no subsidy (see Section 3.1), the total required subsidy is at most $\frac{W-w_{1}}{w_{1}}=\frac{W}{w_{1}}-1$.
We complete the proof of Section 5 by demonstrating that Algorithm 3 runs in polynomial-time. We represent the valuations using a binary matrix $X$ where $v_{i}(o_{j})=1\Longleftrightarrow X(i,j)=1$. Hence, the allocation of items to the bundle of $i_{0}$ at line 1 can be accomplished in $O(mn)$ time.
At each iteration $t$ of the while loop, either $A_{i_{0}}$ or $R$ reduced by 1, ensuring that the loop runs at most $m+n$ times.
@@
According to viswanathan2023general, finding a transfer path starting from agent $i\in N$ (or determining that no such path exists) takes $O(T_{v}\log m)$. Removing agents at the start of each iteration incurs a complexity of $O(nT_{v}\log m)$. Furthermore, as stated in viswanathan2023general, identifying $i^{t}$ requires $O(nT_{v})$. Updating the allocation based on the transfer path, according to the same source, takes $O(m)$.
Thus, each iteration has a total complexity of
$O(nT_{v}\log m+nT_{v}+T_{v}\log m+m)=O\left(nT_{v}\log m+m\right).$
@@
In Appendix A.2, we present a tighter bound that is closer to the lower bound, along with a detailed discussion on its tightness.
Notice that since the output allocation from Algorithm 3 is non-redundant, $A$ maximizes the social welfare. Moreover, as shown in Section 5, $A$ might not be WEF(1,0) (No matter which item is removed from $i_{2}$’s bundle, $i_{1}$ still envies). However, it is WEF(0,1). {proposition} For additive binary valuations, Algorithm 3 computes a $WEF(0,1)$ allocation.
@@
We prove by induction that at the end of each iteration $t\in[T]$ $A^{t}$ satisfies $WEF(0,1)$. This means that for every $i,j\in N$, there exists a set of items $B\subseteq A_{j}^{t}$ of size at most $1$ such that $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t})}{w_{j}}$.
@@
    $v_{i}(A_{j}^{t})=v_{i}(A_{j}^{t-1})$: This is the case where $j$ was not included in a transfer path, or was included but was not the first agent in the path, and exchanged an item for a new one, both having the same value for $i$. By the induction assumption, there exists some singleton $B^{t-1}\subseteq A_{j}^{t-1}$ such that $\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}$. There exists some singleton $B^{t}\subseteq A_{j}^{t}$, with $v_{i}(B^{t})=v_{i}(B^{t-1})$. Hence,
    $\frac{v_{i}(A_{i}^{t})+v_{i}(B^{t})}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B^{t-1})}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}.$
@@
    $v_{i}(A_{j}^{t})<v_{i}(A_{j}^{t-1})$: This is the case where $j$ was included in a transfer path but exchanged an item $i$ values for an item $i$ does not value. Then
    $\frac{v_{i}(A_{i}^{t})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+v_{i}(B)}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})}{w_{j}}>\frac{v_{i}(A_{j}^{t})}{w_{j}}$
    for a set $B\subseteq A_{j}^{t}$ of size at most $1$.
@@
      We can conclude that
      $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}\geq\frac{v_{i}(A_{i}^{t-1})+1}{w_{i}}\geq\frac{v_{i}(A_{j}^{t-1})+1}{w_{j}}=\frac{v_{i}(A_{j}^{t})}{w_{j}}.$
      The claim holds for $B=\{o\}\subseteq A_{j}^{t}$.
@@
      $\frac{v_{i}(A_{i}^{t})+1}{w_{i}}<\frac{v_{i}(A_{j}^{t})}{w_{j}}\leq\frac{v_{j}(A_{j}^{t})}{w_{j}}=\frac{v_{j}(A_{j}^{t^{\prime}})}{w_{j}}=$
      $\frac{v_{j}(A_{j}^{t^{\prime}-1})+1}{w_{j}}\leq\frac{v_{i}(A_{i}^{t^{\prime}-1})+1}{w_{i}}\leq\frac{v_{i}(A_{i}^{t})+1}{w_{i}},$